Subject: APO 3.5m users committee minutes, 10/11/10 meeting
From: Suzanne Hawley
Submitted: Wed, 13 Oct 2010 14:56:44 -0700 (PDT)
Message number: 1206
(previous: 1205,
next: 1207
up: Index)
APO 3.5-m Users Committee Phonecon, 10/11/10
****** Please send science highlights to your users committee
representative, and to Suzanne Hawley and Mark Klaene, by
Friday October 15 for inclusion in the ARC 3.5m Annual Report
We prefer a powerpoint slide including figure and text, but will
accept a figure (jpeg) and text separately. **************
Attending: Suzanne Hawley, Mark Klaene, Michael Strauss, Bill Ketzeback,
Russet McMillan, Remy Indebetouw, Scott Anderson, John Bally, Al Harper,
Sean Moran, Bruce Gillespie
Absent: Jon Holtzman
**********************************
User feedback, comments from institutional representatives:
Chicago (Al Harper) - Al had nothing new to report.
New Mexico State (Jon Holtzman) - no report.
Washington (Scott Anderson) - Scott mentioned that a UW user
recently lost some observing time because of an inability to transfer
observing data. This took about an hour to fix, and Scott wondered if
others are experiencing the same issue. Russet said that there have been
three recent cases, and that it is a new, evolving problem that the APO IT
staff are aware of and are looking for a fix or workaround. Until
recently, firewalls at user institutions assumed a single IP address for
TUI and data downloads. Some firewalls are now expecting multiple IPs for
an observing session, which seems to be causing the problem. Until this
is addressed, users should try to keep taking data if possible during file
transfer blockages, and/or download their data using FTP or other means.
If a quick look at the blocked data is needed, the Observing Specialists
can also help in real time.
Colorado (John Bally) - John had nothing new to report.
Princeton (Michael Strauss) - Michael had nothing new to report.
Johns Hopkins (Sean Moran) - Sean had nothing new to report.
Virginia (Remy Indebetouw) - Remy had nothing new to report.
**********************************
Discussion of telescope/instruments report - Klaene:
Mark highlighted his written report, given below. It is now seasonably
dry, and should hopefully stay like this for the next few weeks. The
winter forecast is for drier than normal conditions, also. Coincident
with tuning the primary mirror position servo, the optical alignment of
the telescope got out of whack; Russet recollimated the telescope but the
root cause is unknown at present. During a recent cloudy night, one of
the UPS units on the telescope failed, which was fixed the next day with a
workaround until we can install replacement parts. The instruments are
generally in good shape, especially after replacing a failed ion pump
controller on SPIcam. Late last week, TUI was upgraded to 1.8.3, which
solved a compatibility problem with Macs using OS 10.4. Suzanne said that
we don't have much of a feeling about how many users are still using this
older operating system. Since the new TUI was upgraded to fix issues with
Agile, users who don't use Agile can still use older versions of TUI.
===> ACTION: Users Committee members should poll their users to find out
how many are using TUI with Mac OS 10.4 and report at the next Users
Committee meeting.
Concerning the engineering plans and CIF projects that are under
consideration (see below), Mark noted that the realuminization of the
primary mirror is slated for next July, and will result in a summer
shutdown that lasts somewhat longer than in the last few years that
haven't included a primary realuminizing. We also intend to begin the
upgrades to the echelle, improve the DIS slit viewer, port the TCC
software to a modern (i.e., not VMS) computer platform, improve dew point
monitoring to both enable more observing close to closing conditions while
avoiding fogging or frosting the optics, reliably enable telescope
tracking up to 85 degrees altitude, complete the implementation of the
top-end metrology (accelerometers and strain gauges), and ramp up research
on telescope enclosure cooling and dome seeing. Upgrades to the
calibration lamps are on hold, pending needed advances in lamp technology.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights, 9/6/10 through 10/05/10
0) Overview
A few days of clear and dry weather, but still more rain and clouds.
Long term forecast is for a winter La Nina so probably a dryer than normal
winter is coming.
1) Telescope
During some PMSS investigation we discovered the telescope was out of
collimation. A significant change to secondary translation corrected the
collimation but we are still investigating what caused it. A failure in
the main UPS unit that powers much of the telescope caused loss of power
around 4:30 AM (during a cloudy night). The failure was traced to a
shorted temperature sensor. A patch was implemented to get the unit back
up quickly and new parts are on order. Minor TUI upgrades were also
released. The problem with Mac OS X 10.4 has been fixed in the most
recent version TUI 1.8.3.
2) Instruments
DIS, TripleSpec, NIC-FPS, the Echelle, and Agile ran with no problems.
SPIcam ION pump controller failed. Upon replacement and pump down,
temperatures are now running normally.
3) Preliminary list of CIF projects for 2011:
Re-aluminize 3.5m
Telescope Control Computer (TCC) Replacement
Echelle Upgrade
DIS Slitviewer mounting improvement and possible replacement
Improve accuracy in dew point metrology
Pointing/Tracking improvements
Top End Metrology
Telescope/Enclosure cooling and dome seeing
**********************************
Science highlights, references for annual report - Hawley:
Suzanne said that she needs science highlights for the ARC 3.5m Annual Report
(she's received nothing so far). They are due this Friday, Oct 15 and should
be sent to the institutional users representative, Suzanne and Mark.
Separately, please send references for publicshed or in press papers,
theses, etc. to Suzanne, Mark and Rene Walterbos who has volunteered again
to collate them (thanks Rene!). Note that we don't include abstracts and
in-preparation papers in our publications list. Suzanne mentioned that we
compare fairly well with similar compilations from other mid-sized
telescopes.
**********************************
Echelle upgrade - Hawley:
Suzanne reported that Jon Holtzman is working on the recommendations for a
new detector, grating, and optical coatings. A new proposal is to modify
the instrument's front end to accommodate a fiber feed and multiple slits.
After some discussion, it appears that recoating the optics may be
infeasible and not worth the effort, especially since it would involve a
difficult realignment of the instrument. The detector upgrade looks most
promising and is presently being explored. Jon is writing a proposal for
presentation at the November ARC Board meeting.
**********************************
APO futures proposals - Hawley:
Suzanne asked if people were working on pre-proposals, which are due 22
October. We need to be proactive now if there are major new initiatives
to be considered, although most of the proposals received will likely be
related to the 2.5-meter telescope. Suzanne discussed some of the ideas
for the 3.5-meter telescope that she was aware of, and there may come a
point where we will ask for expressions of interest for proposals for new
instrumentation, particularly for those coming from non-ARC entities.
**********************************
ACTION ITEMS:
[open from previous meeting]:
===> ACTION: Users Committee members to poll their users to see whether
the current performance of NIC-FPS is adequate for their science.
STATUS: This action is CLOSED. JHU users were asked, but there
was no response; Princeton users are aware of the limitations, and it
impacts some science programs; because of the NIC-FPS noise limitations,
some CU users are making use of NEWFIRM, especially for narrow-band
observations; and nobody seems to be using NIC-FPS at UVa. Mark said that
the plan to try to fix the excessive noise problem involves taking the
instrument to Colorado, likely not until next summer, and to do the tests
and repairs directly on the science, not the engineering, detector -- so
there is some risk that is considered small but significant. With this
said, it was decided to close the action item.
[new actions from this meeting]:
===> ACTION: Users Committee members should poll their users to find out
how many are using TUI with Mac OS 10.4 and report at the next Users
Committee meeting.
**********************************
Next meeting: The next Users Committee phonecon will be on Monday 6
December (skipping November) at 8:30 AM Pacific Time. The agenda and
other materials will be sent to the committee members during the preceding
week.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 1206 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO